A woman deposits $50,000 in a savings account with 4% continuously compounded interest. How many years must she wait until the balance has doubled?
Answer by orca(409)

Suppose her money will be doubled at the end of n years.
At the end of n years she will have 50,000(1 + 4%)^n.
Set it equal to 2*50,000, we have
50,000(1 + 4%)^n = 2*50,000
Dividing both sides by 50,000, we have
(1 + 4%)^n = 2
1.04^n = 2
n = log2/log1.04 = 17.7
So she need to wait 18 years.